# I Don’t Know You, But I Can Catch You: Real-Time Defense against Diverse Adversarial Patches for Object Detectors

### v1

@@ 5.2. Effectiveness
Targeted Hiding Attack. For targeted HA, we use the INRIA dataset (Maggiori et al., 2017), an object detection dataset with annotations for humans, to evaluate the defensive effectiveness against different kinds of patches. For YOLOv2, we use the patches from existing attacks (Thys et al., 2019; Xu et al., 2019; Wu et al., 2020; Hu et al., 2021, 2022a). Specifically, (Thys et al., 2019) provides four types of patches: those minimizing only the objectness score (denoted as (a)), only the classification score (denoted as (b)), and both (denoted as (c) and (d)). For other models with less or even no existing patches, we exploit the code of (Thys et al., 2019) to generate patches, as a complementary to existing attacks. We test the detection accuracy of each patch with and without applying realistic transformations (such as blur, rotation, translation, etc.) to verify their efficiency and their robustness against transformations. The experimental results are shown in Table 1, where each patch has two evaluation results: the white-background result is obtained from evaluating the patch without any transformations, while the gray-background result is obtained from evaluating the patch with transformations.
@@
Impact of different patch sizes. We extend our experiments to investigate the impact of various sizes of AdvPatch (Thys et al., 2019) on the $\text{AP}_{50}$ of YOLOv2 detection on INRIA, as detailed in Table 4. Here, size 0.2 corresponds to the size utilized in Section 5.2. We also evaluated the detection accuracy of the object detector when adversarial patches were replaced with gray patches of the same size. And the ground truth represents the $\text{AP}_{50}$ on the clean dataset. While enlarging the patch size does enhance the attack success rate, the detection performance with NutNet consistently aligns with the performance on gray patches, exhibiting a marginal difference of only approximately 2-3%. This finding suggests that NutNet can effectively detect patches of various sizes.
@@ 5.3. Robustness
where $p$ is the adversarial patch, $(x,y_{gt})$ represents the clean image and its ground-truth annotations, $D$ is the victim object detector, $L$ is the objective function of adversarial attack and $A$ is the patch apply function which applies the patch to the clean image $x$. Note that the attack cannot optimize a patch directly by minimizing $L$ of the reconstructed image $\mathcal{E}(A(p,x))$, since $\mathcal{E}$ consists of reconstruction and dual-mask generation, where the latter is non-differentiable.
We utilize the aforementioned formulation to generate adaptive adversarial patches for hiding persons against the YOLOv2-v4 object detection model. Subsequently, testing is conducted on the INRIA dataset, following the configurations outlined in Section 5.2. Under the condition of a larger $\alpha$, the adaptive patches indeed exhibit a higher evasion rate against NutNet. However, compared to the adversarial patches employed in the previous experiments, the attack performance of these adaptive patches is notably inferior. Conversely, when a smaller $\alpha$ is set, the adaptive patches still maintain a reasonable level of attack effectiveness but remain unable to bypass NutNet. The experimental results are presented in Table 5.
@@
Figure 3. The dot product between the gradient of attacking the object detector and the gradient of bypassing NutNet.
We analyze why adaptive patches cannot simultaneously attack the object detector and bypass NutNet. Using Equation 3, we generate an adaptive adversarial patch, recording adversarial attack loss against the object detector and reconstruction loss against NutNet iteratively. We then calculate and normalize the gradients of the adversarial patch for these two losses, computing their dot product, as shown in Figure 3. In general, we observe that the dot product is consistently negative during the iterative process, indicating an angle consistently greater than $90^{\circ}$ between the two gradients. This implies that the adaptive patch, aiming to reduce the adversarial attack loss, unavoidably increases the reconstruction loss against NutNet, and vice versa. Consequently, it is challenging for this adversarial patch to simultaneously attack the object detector and bypass NutNet.
@@ 5.4. Efficiency
Experimental results indicate that LGS and NutNet have little impact on the inference efficiency of the vanilla model, with only 5.5% and 8.6% overhead respectively. Both LGS and NutNet involve adding another small module for image pre-processing before the detection model, which has a very simple architecture and a very short processing time. Such a small additional overhead hardly affects the real-time detection capability of the detector. The U-Net used in SAC has a large number of parameters (over 1M), leading to a longer forward inference time. As a result, the detection speed of the model under SAC defense is typically reduced by around 50%. Jedi requires the image to be divided into blocks of different sizes, and it calculates the two-dimensional entropy for generating masks. Additionally, Jedi uses an autoencoder to cluster discrete masks, resulting in a significant computational overhead. For OS, however, its sliding mask and unionizing boxes operations are highly time-consuming, resulting in significant processing time for detecting a single image, often taking several seconds. In conclusion, LGS and NutNet are capable of meeting the real-time requirements of defending object detectors, while OS is too time-consuming and cannot fulfill real-time demands.
@@ 5.5. Comparison with Existing Defenses
SAC provides a slight enhancement in detection accuracy for clean images (Table 2). However, its effectiveness against adversarial patches is limited, and the defensive outcome varies (Table 1 and 3). Jedi demonstrates slight defensive capabilities against AA attacks (Table 3), but its defensive effectiveness is inconsistent for HA attacks (Table 1 and 2). Sometimes it provides a defense, while sometimes it may have a negative impact, challenging Jedi’s assumption that adversarial patch entropy surpasses that of clean images. OS defends against adversarial patches in corners using sliding masks but shows mixed results. It significantly impacts YOLOv2, SSD, and Faster RCNN but performs better in YOLOv4 (Table 1). The method’s high computational cost, primarily due to sliding masks, raises concerns about real-time detection (Figure 5). LGS has a certain level of defense against various types of attacks (Table 1 and 2). However, its effectiveness in defending against AA is unstable, sometimes exhibiting defense capabilities and at other times further enhancing the attack performance (Table 3).
@@ 5.6. Effectiveness in the Physical World
We also conduct tests on the effectiveness of defense against HA and AA in the physical world. Demos can be found at: https://sites.google.com/view/nutnet. The goal of untargeted HA is to hide all objects in the scene. However, measuring the success rate of the attack based on this criterion is not ideal because not all objects can be consistently hidden. Therefore, we measure the attack success rate by evaluating whether it can hide the person in the scene, similar to targeted HA. Specifically, we print out different types of adversarial patches and carry out physical attacks. We record videos of the camera and the object performing relative motion and input them into the detector. Each patch’s video lasted for more than 40 seconds (30 FPS). We calculate the success rate of the adversarial attacks in each video, where the success rate is defined as the percentage of frames in which the attack succeeded out of the total frames. Then, we apply NutNet to process the videos before inputting them into the target detector, and we calculate the attack success rate in each video. Each patch is evaluated on the model it is specifically targeted against, rather than employing transferability attacks. The experimental results are shown in Table 7. Regardless of the type of attack, NutNet is able to significantly reduce the attack success rate to a very low level. Note that the original attack success rate of NaturalPatch(Hu et al., 2021) is already low, which is consistent with the evaluation in the digital world in Section 5.2. More physical evaluations will be demonstrated in Appendix C.1.
@@ 5.7. Ablation Study
Impact of the Threshold $\kappa$. The threshold $\kappa$ affects NutNet’s sensitivity to suspicious blocks. Therefore, we also tested the detection accuracy of YOLOv2 with different NutNet thresholds, as shown in Table 10. Overall, as $\kappa_{1}$ and $\kappa_{2}$ increase, the detection accuracy on both clean images and images with HA patches improves, likely due to the larger thresholds resulting in a very low false positive rate for NutNet. However, larger $\kappa_{1}$ and $\kappa_{2}$ lead to poorer defense against AA patches, as the masking rate for these patches decreases, causing more false negatives. In summary, selecting appropriate thresholds is crucial for balancing the false positive and false negative rates effectively.
@@ 6. Discussion
Table 12. Performance and efficiency of NutNet-13 with and without image inpainting under different patches against YOLOv2 detector.

### v2

@@ 5.2. Effectiveness
Targeted Hiding Attack. For targeted HA, we use the INRIA dataset (Maggiori et al., 2017), an object detection dataset with annotations for humans, to evaluate the defensive effectiveness against different kinds of patches. For YOLOv2, we use the patches from existing attacks (Thys et al., 2019; Xu et al., 2019; Wu et al., 2020; Hu et al., 2021, 2022a). Specifically, (Thys et al., 2019) provides four types of patches: those minimizing only the objectness score (denoted as (a) in Table 1), only the classification score (denoted as (b)), and both (denoted as (c) and (d)). For other models with less or even no existing patches, we exploit the code of (Thys et al., 2019) to generate patches, as a complementary to existing attacks. We test the detection accuracy of each patch with and without applying realistic transformations (such as blur, rotation, translation, etc.) to verify their efficiency and their robustness against transformations. The experimental results are shown in Table 1, where each patch has two evaluation results: the white-background result is obtained from evaluating the patch without any transformations, while the gray-background result is obtained from evaluating the patch with transformations.
@@
Impact of different patch sizes. We extend our experiments to investigate the impact of various sizes of AdvPatch (Thys et al., 2019) on the $\text{AP}_{50}$ of YOLOv2 detection on INRIA, as detailed in Table 4. Here, size 0.2 corresponds to the size utilized in Section 5.2. We also evaluate the detection accuracy of the object detector when adversarial patches are replaced with gray patches of the same size. And the ground truth represents the $\text{AP}_{50}$ on the clean dataset. While enlarging the patch size does enhance the attack success rate, the detection performance with NutNet consistently aligns with the performance on gray patches, exhibiting a marginal difference of only approximately 2-3%. This finding suggests that NutNet can effectively detect patches of various sizes.
@@ 5.3. Robustness
where $p$ is the adversarial patch, $(x,y_{gt})$ represents the clean image and its ground-truth annotations, $D$ is the victim object detector, $L$ is the objective function of adversarial attack and $A$ is the patch apply function which applies the patch to the clean image $x$. Note that the attacker cannot optimize a patch directly by minimizing $L$ of the reconstructed image $\mathcal{E}(A(p,x))$, since $\mathcal{E}$ consists of reconstruction and dual-mask generation, where the latter is non-differentiable.
We utilize the aforementioned formulation to generate adaptive adversarial patches for hiding persons against the YOLOv2-v4 detectors. Subsequently, testing is conducted on the INRIA dataset, following the configurations outlined in Section 5.2. Under the condition of a larger $\alpha$, the adaptive patches indeed exhibit a higher evasion rate against NutNet. However, compared to the adversarial patches employed in the previous experiments, the attack performance of these adaptive patches is notably inferior. Conversely, when a smaller $\alpha$ is set, the adaptive patches still maintain a reasonable level of attack effectiveness but remain unable to bypass NutNet. The experimental results are presented in Table 5.
@@
Figure 3. The inner product between the gradient of attacking the object detector and the gradient of bypassing NutNet.
We analyze why adaptive patches cannot simultaneously attack the object detector and bypass NutNet. Using Equation 3, we generate an adaptive adversarial patch, recording adversarial attack loss against the object detector and reconstruction loss against NutNet iteratively. We then calculate and normalize the gradients of the adversarial patch for these two losses, computing their inner product, as shown in Figure 3. In general, we observe that the inner product is consistently negative during the iterative process, indicating an angle consistently greater than $90^{\circ}$ between the two gradients. This implies that the adaptive patch, aiming to reduce the adversarial attack loss, unavoidably increases the reconstruction loss against NutNet, and vice versa. Consequently, it is challenging for this adversarial patch to simultaneously attack the object detector and bypass NutNet.
@@ 5.4. Efficiency
Experimental results indicate that LGS and NutNet have little impact on the inference efficiency of the vanilla model, with only 5.5% and 8.6% overhead respectively. Both LGS and NutNet involve adding another small module for image pre-processing before the detection model, which has a very simple architecture and a very short processing time. Such a small additional overhead hardly affects the real-time detection capability of the detector. The U-Net used in SAC has a large number of parameters (over 1M), leading to a longer forward inference time. As a result, the detection speed of the model under SAC defense is typically reduced by around 50%. Jedi requires the image to be divided into blocks of different sizes, and it calculates the two-dimensional entropy for generating masks. Additionally, Jedi uses an autoencoder to cluster discrete masks, resulting in a significant computational overhead. For OS, however, its sliding mask and unionizing boxes operations are highly time-consuming, resulting in significant processing time for detecting a single image, often taking several seconds. In conclusion, LGS and NutNet are capable of meeting the real-time requirements of defending object detectors, while the others are too time-consuming and cannot fulfill real-time demands.
@@ 5.5. Comparison with Existing Defenses
SAC provides a slight enhancement in detection accuracy for clean images (Table 2). However, its effectiveness against adversarial patches is limited, and the defensive outcome varies (Table 1 and 3). Jedi demonstrates slight defensive capabilities against AA (Table 3), but its defensive effectiveness is inconsistent for HA (Table 1 and 2). Sometimes it provides a defense, while sometimes it may have a negative impact, challenging Jedi’s assumption that adversarial patch entropy surpasses that of clean images. OS defends against adversarial patches in corners using sliding masks but shows mixed results. It significantly impacts YOLOv2, SSD, and Faster RCNN but performs better in YOLOv4 (Table 1). The method’s high computational cost, primarily due to sliding masks, raises concerns about real-time detection (Figure 5). LGS has a certain level of defense against various types of attacks (Table 1 and 2). However, its effectiveness in defending against AA is unstable, sometimes exhibiting defense capabilities and at other times further enhancing the attack performance (Table 3).
@@ 5.6. Effectiveness in the Physical World
We also conduct tests on the effectiveness of defense against HA and AA in the physical world. Demos can be found at: https://sites.google.com/view/nutnet. The goal of untargeted HA is to hide all objects in the scene. However, measuring the success rate of the attack based on this criterion is not ideal because not all objects can be consistently hidden. Therefore, we measure the attack success rate by evaluating whether it can hide the person in the scene, similar to targeted HA. Specifically, we print out different types of adversarial patches and carry out physical attacks. We record videos of the camera and the object performing relative motion and input them into the detector. Each patch’s video lasted for more than 40 seconds (30 FPS). We calculate the success rate of the adversarial attacks in each video, where the success rate is defined as the percentage of frames in which the attack succeeded out of the total frames. Then, we apply NutNet to process the videos before inputting them into the detector, and we calculate the attack success rate in each video again. Each patch is evaluated on the model it is specifically targeted against, rather than employing transferability attacks. The experimental results are shown in Table 7. Regardless of the type of attack, NutNet is able to significantly reduce the attack success rate to a very low level. Note that the original attack success rate of NaturalPatch(Hu et al., 2021) is already low, which is consistent with the evaluation in the digital world in Section 5.2. More physical evaluations will be demonstrated in Appendix C.1.
@@ 5.7. Ablation Study
Impact of the Threshold $\kappa$. The threshold $\kappa$ affects NutNet’s sensitivity to suspicious blocks. Therefore, we also tested the detection accuracy of YOLOv2 with different NutNet thresholds, as shown in Table 10. Overall, as $\kappa_{1}$ and $\kappa_{2}$ increase, the detection accuracy on both clean images and images with HA patches improves, likely due to the larger thresholds resulting in a very low false positive rate for NutNet. However, larger $\kappa_{1}$ and $\kappa_{2}$ lead to poorer defense against AA patches, as the masking rate for these patches decreases, causing more false negatives. In summary, selecting appropriate thresholds is crucial for balancing the false positive of defending against HA and false negative of defending against AA effectively.
@@ 6. Discussion
Table 12. Performance and efficiency of NutNet-13 with and without image inpainting under different patches against YOLOv2 detector on INRIA.